# Self-Paced and Self-Corrective Masked Prediction for Movie Trailer Generation

### v1

@@ 1 Introduction
Movie trailers play a vital role in the movie industry, serving as a key medium for attracting audiences and shaping their first impressions of a movie. Typically, trailer editing involves iteratively selecting and arranging shot sequences to ensure coherence and harmony throughout the generated trailer, making it an exceptionally challenging and time-consuming task that requires semantic video understanding. To make the entire process more efficient, numerous efforts [19, 34, 7, 39, 40, 1, 35, 54] have been made to automate the generation of movie trailers.
@@
Early methods [19, 34, 7, 39, 40, 35] are mostly rule-based, relying on additional manually-designed editing rules and thus suffering from poor generalizability. With the development of deep learning, several data-driven methods have been proposed [50, 1, 54, 44, 21] to directly learn trailer composition patterns from paired movie–trailer data. In principle, these methods yield the same “selection-then-ranking” paradigm shown in Fig. 1(a), which decomposes trailer generation into two steps: shot selection and ranking. This paradigm prevents the model from jointly reasoning about the semantic relevance and temporal continuity among shots. Recently, some attempts have been made to achieve selection and ranking jointly, formulating the task as a sequence-to-sequence problem [1, 21]. As shown in Fig. 1(b), these methods predict trailer shots conditioned on the given movie in an auto-regressive manner. However, both selection-then-ranking and auto-regression lack self-correction mechanisms to revise early shot selections/predictions, which fundamentally differ from how human editors work, as professional editors repeatedly refine shot connections during the trailer editing process. As a result, these two paradigms suffer from inevitable error propagation, which limits the quality of the generated trailers.
@@
To address the above issues, we formulate the trailer generation task as a masked prediction problem and propose a novel Self-paced and Self-corrective Masked Prediction (SSMP) method. In particular, SSMP learns a Transformer encoder-based trailer generator with a self-correction mechanism in a self-paced learning framework. As shown in Fig. 1(c), the model takes the given movie shots as the prompt and reconstructs the masked trailer shots iteratively. At each step, the model predicts shots of all masked positions simultaneously and re-masks low-confidence shots for subsequent iterations. This generation process effectively alleviates the error propagation issue — it considers shot selection and ranking jointly and incorporates a self-correction mechanism that revises the generated shots. When training our model, we are inspired by self-paced learning strategies [24, 20], designing a self-paced mask ratio scheduler to improve training efficiency, stability, and model performance. The scheduler dynamically adjusts the mask ratio of the shot sequences according to the capability of the model in the training process.
@@ 2.1 Automatic Trailer Generation
As aforementioned, early automatic trailer generation methods are rule-based, leveraging meta or side information to summarize common trailer grammar or editing templates. These methods can extract key components (e.g., logo, theme music, and so on) of trailers from movies [19, 22]. Some of them further identify turning points and cinematic grammars from movies for guiding trailer generation [34, 7], which leads to useful editing rules for specific movie genres [39, 40, 35]. Such rule-based methods suffer from poor generalization, since the predefined rules cannot cover various movie narratives and visual styles.
To address these issues, data-driven methods are proposed to learn trailer generators from movie-trailer pairs in supervised [37] or unsupervised [30, 47] manner. These methods commonly apply the “selection-then-ranking” paradigm, selecting candidate shots for trailers and arrange them accordingly. Among them, various measurements are used as shot selection criteria, including the visual attractiveness [51], the shot expressiveness [38], the shot emotional intensity [18], and the shot relevance based on side information (e.g., viewer preference [25], subtitles [17], movie summaries [12, 3, 2, 23]). For shot ranking, most existing methods rely on external guidance [26], such as music-video semantic consistency [44, 50, 29, 54] and narrative assistance from large language models [3, 2, 27].
However, the selection-then-ranking paradigm suffers from inherent limitations, as it decouples shot selection and ranking despite their strong interdependence. To unify these two steps, the work in [1] draws inspiration from machine translation and formulates the task as a sequence-to-sequence prediction. Building upon this work, the subsequent study [21] incorporates the movie’s main plot as conditional inputs, and further generates an accompanying soundtrack based on the generated trailer. Similar to existing language models, these two methods apply an auto-regressive (AR) paradigm, predicting each shot based on the given movie shots and the previous predictions. However, without any self-correction mechanism, these AR-based methods are still different from how human editors work — human editors usually polish a trailer iteratively, repeatedly selecting and replacing shots across different shot positions instead of following a strictly sequential order.
@@ 2.2 Masked Prediction for Generation
As a powerful learning paradigm, masked prediction has been widely used in pre-training foundation models. Typically, BERT [9] achieves a pre-trained large-scale language model by masked prediction, obtaining remarkable success across various natural language processing (NLP) tasks. Motivated by BERT, CMLM [14] achieves conditional masked language modeling, and LLaDA [33] proposes a discrete diffusion model with a random mask ratio, allowing for contextually aware text generation. The success in NLP motivates studies to explore masked prediction for image representation learning, which is often implemented by reconstructing masked pixels [16] and patches [4, 10, 6]. For video representation learning, the BEVT in [48] and the joint visual-linguistic models in [43, 52] are learned to reconstruct masked video tokens, both demonstrating the effectiveness of masked prediction. However, the current methods seldom investigate the impact of mask ratio on the model performance, and none of them consider constructing trailer generators via masked prediction. Our work fills this blank, learning a strong movie trailer generator via a self-paced masked prediction method.
@@ 3 Proposed Method
In this study, we denote a movie and its corresponding trailer as $\mathcal{M}$ and $\mathcal{V}$, respectively. Referring to previous methods [1, 54], the movie and its trailer are segmented by the shot-boundary detector TransNet-v2 [41], leading to two shot sequences, i.e., $\mathcal{M}=\{m_{i}\}_{i=1}^{I}$ and $\mathcal{V}=\{v_{j}\}_{j=1}^{J}$. Typically, each trailer shot is sampled from the movie shots, i.e., $\mathcal{V}\subset\mathcal{M}$. In this work, we formulate the movie trailer generation task as a masked prediction problem conditioned on the given movie. As illustrated in Fig. 3, our model takes the movie shot sequence $\mathcal{M}$ and trailer shot sequence $\mathcal{V}$ as inputs, in which each shot is treated as a token. The model captures both the underlying shot selection mechanism from the source movie and the sequential dependencies among the trailer shots through the reconstruction of masked trailer shots. When generating a trailer, the well-trained model takes the given movie shot sequence as a prompt and progressively fills the trailer sequence from all mask placeholders. During this process, a self-correction mechanism is implemented as re-masking low-confidence shots iteratively.
@@ 3.1 A Masked Prediction Framework for Training
Inspired by BERT [9] and LLaDA [33], our model is trained under a conditional masked prediction framework, where partial input tokens are masked through a forward process and reconstructed through a reverse process. As shown in Fig. 2(a), we utilize a pretrained model ImageBind [15] as our video encoder, extracting movie shot features $\bm{M}=[\bm{m}_{i}]\in\mathbb{R}^{I\times D}$ (i.e., the prompt) and trailer shot features $\bm{V}=[\bm{v}_{j}]\in\mathbb{R}^{J\times D}$ (i.e., the target), respectively, where $D$ denotes the feature dimension. Each trailer shot is considered to originate from the movie shot that has the highest similarity score. The corresponding ground truth labels are formulated as a matrix $\bm{G}=[g_{j,i}]\in\{0,1\}^{J\times I}$, i.e., for j=1,,Jj=1,...,J and i=1,,Ii=1,...,I, we have $s_{j,i}=\frac{\langle\bm{v}_{j},\bm{m}_{i}\rangle}{\|\bm{v}_{j}|\|\bm{m}_{i}\|}$ and
@@
We concatenate $\bm{M}$ and $\bm{V}$ into a single sequence and feed the sequence into the proposed mask predictor $f_{\theta}$, which is implemented as a Transformer encoder [46]. Within this predictor, the prompt part is kept unchanged, while the target features are independently masked at the same ratio, resulting in a partially masked sequence $\bm{V}^{t}=[\bm{v}_{j}^{t}]\in\mathbb{R}^{J\times D}$, i.e., for j=1,,Jj=1,...,J,
@@
Unlike the select-and-ranking methods [51, 50, 54] and the autoregressive ones [1, 21], our model achieves bi-directional contextual modeling of movie trailer. The model is optimized using the AdamW optimizer [31], performing iterative updates until convergence.
@@ 3.2 A Self-Paced Mask Ratio Scheduler
The mask ratio $t$ is critical for the model training, as it determines how much information is hidden from the model and thus reflects the reconstruction difficulty. A low mask ratio results in a trivial task inadequate for model training, while a high mask ratio presents an overly difficult task that prevents the model from converging. Inspired by the self-paced learning strategies [24, 20], we design a self-paced masking strategy that adaptively schedules the mask ratio (i.e., task difficulty) to improve learning efficiency.
@@ 3.3 Trailer Generation with Self-Correction
Given the trained model $f_{\theta}$, we design a new trailer generation process with a progressive self-correction mechanism. Given a movie, we segment it into shots [41] and extract its shot features $\bm{M}\in\mathbb{R}^{I\times D}$ by the ImageBind-based video encoder [15]. To generate a trailer with $J$ shots, [^1] Similar to the work in [54], we determine the number of trailer shots based on the music track associated with the trailer. We set $J$ to match the number of music segments detected by Ruptures [45]. we initialize a shot sequence $\bm{M};\bm{V}_{0}]$, where $\bm{V}_{0}=[\text{MP}]\in\mathbb{R}^{J\times D}$ is a fully masked sequence, and predict the masked features by $f_{\theta}$. In addition, we introduce a vector $\bm{q}=[q_{j}]\in[0,1]^{J}$, in which $q_{j}$ indicates whether a masked position should be filled. This vector is initialized as a zero vector. In principle, our model achieves assigning $J$ distinct movie shots to the masked positions iteratively.
@@ 3.4 Post-Processing for Trailer Generation
Based on the trailer index vector $\bm{z}$ obtained by Algorithm 1, we select the corresponding shots from the given movie and generate a trailer, denoted as $\mathcal{T}_{\bm{z}}$. Following the work in [54], we adopt the music track from the official trailer of the movie for $\mathcal{T}_{\bm{z}}$ and perform the corresponding post-processing steps. We first segment the music track by Ruptures [45] and align the duration of each trailer shot with its corresponding music shot. We then employ DeepSeek-V3 [8] to analyze and select key subtitles as trailer narrations and apply MiniCPM-V2.6 [53] to generate one-sentence descriptions for each trailer shot. To determine the shot positions of the narrations, we utilize CLIP [36] to extract textual representations of narrations and shot descriptions and align them by maximizing their textual similarity using dynamic programming [5]. Please refer to the work in [54] or our supplementary file for more details.
@@
VASNet [11] 0.0712 0.0645 0.0676 100.62 0.43 0.0496 0.0422 0.0455 84.79 0.44
Video CLIP-It [32] 0.0711 0.0629 0.0667 101.12 0.45 0.0832 0.0710 0.0764 85.82 0.38
Summarization OTVS [49] 0.0688 0.0613 0.0648 101.25 0.46 0.0834 0.0711 0.0766 84.82 0.39
Muvee [13] 0.2400 0.0461 0.0714 103.50 0.36
M2T [37] 0.0611 0.0503 0.0515 95.67 0.42
V2T [19] 0.1121 0.0603 0.0945 103.75 0.52
Selection- PPBVAM [51] 0.0813 0.1244 0.0945 101.50 0.53
then-Ranking IPOT [50] 0.1218 0.1425 0.1311 101.50 0.42 0.1258 0.1483 0.1357 83.21 0.46
MMSC [54] 0.1301 0.1449 0.1391 99.25 0.58 0.1852 0.2164 0.1991 82.48 0.50
Autoregression TGT [1] 0.1185 0.1142 0.1153 103.87 0.48 0.1243 0.1433 0.1326 88.29 0.43
@@ 4.1 Implementation Details
Datasets. We extend the CMTD dataset [50] to train our model and evaluate the model on the two testing sets of the dataset, called Test-8 and Test-74, respectively. This dataset contains 500 movies and 922 corresponding trailers. We enrich the dataset by 200 additional recent movies and their corresponding trailers. Among them, 30 movie-trailer pairs newly released in 2024 are selected to construct a new test set, called Test-2024, which is used to further evaluate the model’s generalization ability. The collection of all the three testing sets is called Test-ALL, which is applied in ablation studies. For each video (a movie or a trailer), we segment shots and extract their features following the pre-processing pipeline in [50, 44].
Baselines. We compare our SSMP method with state-of-the-art trailer generation methods, including selection-then-ranking methods (e.g., V2T [19], M2T [37], PPBVAM [51], IPOT [50], and MMSC [54]) and the autoregressive method TGT [^2] Since the data and code are not publicly available, we reproduce TGT [1] by ourselves. When generating a trailer, we set the same length for TGT and our method for a fair comparison. [1, 21]. Here, IPOT [50] and MMSC [54] require background music as input, while the others only leverage visual information. Several video summarization methods are considered when evaluating the shot selection performance, including Muvee [13], VASNet [11], CLIP-It [32], and OTVS [49]. For a fair comparison, all the methods are tested on the same dataset, and all the trailer generation methods apply the post-processing shown in Sec. 3.4.
Evaluation metrics. To evaluate the shot selection performance, we employ Precision, Recall, and F1-score (F1) by comparing each $\bm{z}$ with the ground-truth movie–trailer alignment. Following the work in [28, 1, 21], we further use Levenshtein Distance (LD) and Pairwise Agreement Accuracy (AA) to measure the shot ranking performance. For each generated trailer and the official one, LD quantifies their dissimilarity by computing the minimum number of edits required to convert one shot sequence into another, and AA measures the proportion of well-aligned shot pairs.
@@
VASNet [11] 0.0990 0.0806 0.0886 93.84 0.44
CLIP-It [32] 0.0443 0.0370 0.0403 95.51 0.31
OTVS [49] 0.1021 0.0831 0.0914 93.56 0.44
IPOT [50] 0.1236 0.1460 0.1336 93.82 0.42
MMSC [54] 0.1567 0.1485 0.1506 93.91 0.52
TGT [1] 0.1521 0.1601 0.1601 103.25 0.46

### v2

@@ 1 Introduction
Movie trailers play a vital role in the movie industry, serving as a key medium for attracting audiences and shaping their first impressions of a movie. Typically, trailer editing involves iteratively selecting and arranging shot sequences to ensure coherence and harmony throughout the generated trailer, making it an exceptionally challenging and time-consuming task that requires semantic video understanding. To make the entire process more efficient, numerous efforts [irie2010automatic, papalampidi2023finding, chen2004action, smeaton2006automatically, smith2017harnessing, argaw2024towards, praveen2025video, zhu2025weakly] have been made to automate the generation of movie trailers.
@@
Early methods [irie2010automatic, papalampidi2023finding, chen2004action, smeaton2006automatically, smith2017harnessing, praveen2025video] are mostly rule-based, relying on additional manually-designed editing rules and thus suffering from poor generalizability. With the development of deep learning, several data-driven methods have been proposed [wang2024inverse, argaw2024towards, zhu2025weakly, sun2025multi, jin2025tvmtrailer] to directly learn trailer composition patterns from paired movie–trailer data. In principle, these methods yield the same “selection-then-ranking” paradigm shown in Fig. 1(a), which decomposes trailer generation into two steps: shot selection and ranking. This paradigm prevents the model from jointly reasoning about the semantic relevance and temporal continuity among shots. Recently, some attempts have been made to achieve selection and ranking jointly, formulating the task as a sequence-to-sequence problem [argaw2024towards, jin2025tvmtrailer]. As shown in Fig. 1(b), these methods predict trailer shots conditioned on the given movie in an auto-regressive manner. However, both selection-then-ranking and auto-regression lack self-correction mechanisms to revise early shot selections/predictions, which fundamentally differ from how human editors work, as professional editors repeatedly refine shot connections during the trailer editing process. As a result, these two paradigms suffer from inevitable error propagation, which limits the quality of the generated trailers.
@@
To address the above issues, we formulate the trailer generation task as a masked prediction problem and propose a novel Self-paced and Self-corrective Masked Prediction (SSMP) method. In particular, SSMP learns a Transformer encoder-based trailer generator with a self-correction mechanism in a self-paced learning framework. As shown in Fig. 1(c), the model takes the given movie shots as the prompt and reconstructs the masked trailer shots iteratively. At each step, the model predicts shots of all masked positions simultaneously and re-masks low-confidence shots for subsequent iterations. This generation process effectively alleviates the error propagation issue — it considers shot selection and ranking jointly and incorporates a self-correction mechanism that revises the generated shots. When training our model, we are inspired by self-paced learning strategies [NIPS2010_e57c6b95, jiang2014self], designing a self-paced mask ratio scheduler to improve training efficiency, stability, and model performance. The scheduler dynamically adjusts the mask ratio of the shot sequences according to the capability of the model in the training process.
@@ 2.1 Automatic Trailer Generation
As aforementioned, early automatic trailer generation methods are rule-based, leveraging meta or side information to summarize common trailer grammar or editing templates. These methods can extract key components (e.g., logo, theme music, and so on) of trailers from movies [irie2010automatic, kakimoto2018extraction]. Some of them further identify turning points and cinematic grammars from movies for guiding trailer generation [papalampidi2023finding, chen2004action], which leads to useful editing rules for specific movie genres [smeaton2006automatically, smith2017harnessing, praveen2025video]. Such rule-based methods suffer from poor generalization, since the predefined rules cannot cover various movie narratives and visual styles.
To address these issues, data-driven methods are proposed to learn trailer generators from movie-trailer pairs in supervised [rehusevych2019movie2trailer] or unsupervised [liu2015semi, wang2020learning] manner. These methods commonly apply the “selection-then-ranking” paradigm, selecting candidate shots for trailers and arrange them accordingly. Among them, various measurements are used as shot selection criteria, including the visual attractiveness [xu2015trailer], the shot expressiveness [shaukat2024genvis], the shot emotional intensity [hu2024decoupled], and the shot relevance based on side information (e.g., viewer preference [lee2025hippo], subtitles [hesham2018smart], movie summaries [gaikwad2021plots, balestri2024trailer, balestri2024automatic, kawai2007automated]). For shot ranking, most existing methods rely on external guidance [li2025shot], such as music-video semantic consistency [sun2025multi, wang2024inverse, JCST-2212-13064, zhu2025weakly] and narrative assistance from large language models [balestri2024trailer, balestri2024automatic, li2025shots].
However, the selection-then-ranking paradigm suffers from inherent limitations, as it decouples shot selection and ranking despite their strong interdependence. To unify these two steps, the work in [argaw2024towards] draws inspiration from machine translation and formulates the task as a sequence-to-sequence prediction. Building upon this work, the subsequent study [jin2025tvmtrailer] incorporates the movie’s main plot as conditional inputs, and further generates an accompanying soundtrack based on the generated trailer. Similar to existing language models, these two methods apply an auto-regressive (AR) paradigm, predicting each shot based on the given movie shots and the previous predictions. However, without any self-correction mechanism, these AR-based methods are still different from how human editors work — human editors usually polish a trailer iteratively, repeatedly selecting and replacing shots across different shot positions instead of following a strictly sequential order.
@@ 2.2 Masked Prediction for Generation
As a powerful learning paradigm, masked prediction has been widely used in pre-training foundation models. Typically, BERT [devlin2019bert] achieves a pre-trained large-scale language model by masked prediction, obtaining remarkable success across various natural language processing (NLP) tasks. Motivated by BERT, CMLM [ghazvininejad2019mask] achieves conditional masked language modeling, and LLaDA [nie2025large] proposes a discrete diffusion model with a random mask ratio, allowing for contextually aware text generation. The success in NLP motivates studies to explore masked prediction for image representation learning, which is often implemented by reconstructing masked pixels [he2022masked] and patches [bao2021beit, dong2023peco, chang2022maskgit]. For video representation learning, the BEVT in [wang2022bevt] and the joint visual-linguistic models in [sun2019videobert, yang2020bert] are learned to reconstruct masked video tokens, both demonstrating the effectiveness of masked prediction. However, the current methods seldom investigate the impact of mask ratio on the model performance, and none of them consider constructing trailer generators via masked prediction. Our work fills this blank, learning a strong movie trailer generator via a self-paced masked prediction method.
@@ 3 Proposed Method
In this study, we denote a movie and its corresponding trailer as $\mathcal{M}$ and $\mathcal{V}$, respectively. Referring to previous methods [argaw2024towards, zhu2025weakly], the movie and its trailer are segmented by the shot-boundary detector TransNet-v2 [soucek2024transnet], leading to two shot sequences, i.e., $\mathcal{M}=\{m_{i}\}_{i=1}^{I}$ and $\mathcal{V}=\{v_{j}\}_{j=1}^{J}$. Typically, each trailer shot is sampled from the movie shots, i.e., $\mathcal{V}\subset\mathcal{M}$. In this work, we formulate the movie trailer generation task as a masked prediction problem conditioned on the given movie. As illustrated in Fig. 3, our model takes the movie shot sequence $\mathcal{M}$ and trailer shot sequence $\mathcal{V}$ as inputs, in which each shot is treated as a token. The model captures both the underlying shot selection mechanism from the source movie and the sequential dependencies among the trailer shots through the reconstruction of masked trailer shots. When generating a trailer, the well-trained model takes the given movie shot sequence as a prompt and progressively fills the trailer sequence from all mask placeholders. During this process, a self-correction mechanism is implemented as re-masking low-confidence shots iteratively.
@@ 3.1 A Masked Prediction Framework for Training
Inspired by BERT [devlin2019bert] and LLaDA [nie2025large], our model is trained under a conditional masked prediction framework, where partial input tokens are masked through a forward process and reconstructed through a reverse process. As shown in Fig. 2(a), we utilize a pretrained model ImageBind [girdhar2023imagebind] as our video encoder, extracting movie shot features $\bm{M}=[\bm{m}_{i}]\in\mathbb{R}^{I\times D}$ (i.e., the prompt) and trailer shot features $\bm{V}=[\bm{v}_{j}]\in\mathbb{R}^{J\times D}$ (i.e., the target), respectively, where $D$ denotes the feature dimension. Each trailer shot is considered to originate from the movie shot that has the highest similarity score. The corresponding ground truth labels are formulated as a matrix $\bm{G}=[g_{j,i}]\in\{0,1\}^{J\times I}$, i.e., for j=1,,Jj=1,...,J and i=1,,Ii=1,...,I, we have $s_{j,i}=\frac{\langle\bm{v}_{j},\bm{m}_{i}\rangle}{\|\bm{v}_{j}|\|\bm{m}_{i}\|}$ and
@@
We concatenate $\bm{M}$ and $\bm{V}$ into a single sequence and feed the sequence into the proposed mask predictor $f_{\theta}$, which is implemented as a Transformer encoder [vaswani2017attention]. Within this predictor, the prompt part is kept unchanged, while the target features are independently masked at the same ratio, resulting in a partially masked sequence $\bm{V}^{t}=[\bm{v}_{j}^{t}]\in\mathbb{R}^{J\times D}$, i.e., for j=1,,Jj=1,...,J,
@@
Unlike the select-and-ranking methods [xu2015trailer, wang2024inverse, zhu2025weakly] and the autoregressive ones [argaw2024towards, jin2025tvmtrailer], our model achieves bi-directional contextual modeling of movie trailer. The model is optimized using the AdamW optimizer [loshchilov2017decoupled], performing iterative updates until convergence.
@@ 3.2 A Self-Paced Mask Ratio Scheduler
The mask ratio $t$ is critical for the model training, as it determines how much information is hidden from the model and thus reflects the reconstruction difficulty. A low mask ratio results in a trivial task inadequate for model training, while a high mask ratio presents an overly difficult task that prevents the model from converging. Inspired by the self-paced learning strategies [NIPS2010_e57c6b95, jiang2014self], we design a self-paced masking strategy that adaptively schedules the mask ratio (i.e., task difficulty) to improve learning efficiency.
@@ 3.3 Trailer Generation with Self-Correction
Given the trained model $f_{\theta}$, we design a new trailer generation process with a progressive self-correction mechanism. Given a movie, we segment it into shots [soucek2024transnet] and extract its shot features $\bm{M}\in\mathbb{R}^{I\times D}$ by the ImageBind-based video encoder [girdhar2023imagebind]. To generate a trailer with $J$ shots, [^1] Similar to the work in [zhu2025weakly], we determine the number of trailer shots based on the music track associated with the trailer. We set $J$ to match the number of music segments detected by Ruptures [truong2020selective]. we initialize a shot sequence $\bm{M};\bm{V}_{0}]$, where $\bm{V}_{0}=[\text{MP}]\in\mathbb{R}^{J\times D}$ is a fully masked sequence, and predict the masked features by $f_{\theta}$. In addition, we introduce a vector $\bm{q}=[q_{j}]\in[0,1]^{J}$, in which $q_{j}$ indicates whether a masked position should be filled. This vector is initialized as a zero vector. In principle, our model achieves assigning $J$ distinct movie shots to the masked positions iteratively.
@@ 3.4 Post-Processing for Trailer Generation
Based on the trailer index vector $\bm{z}$ obtained by Algorithm 1, we select the corresponding shots from the given movie and generate a trailer, denoted as $\mathcal{T}_{\bm{z}}$. Following the work in [zhu2025weakly], we adopt the music track from the official trailer of the movie for $\mathcal{T}_{\bm{z}}$ and perform the corresponding post-processing steps. We first segment the music track by Ruptures [truong2020selective] and align the duration of each trailer shot with its corresponding music shot. We then employ DeepSeek-V3 [deepseekai2024deepseekv3technicalreport] to analyze and select key subtitles as trailer narrations and apply MiniCPM-V2.6 [yao2024minicpm] to generate one-sentence descriptions for each trailer shot. To determine the shot positions of the narrations, we utilize CLIP [radford2021learning] to extract textual representations of narrations and shot descriptions and align them by maximizing their textual similarity using dynamic programming [bellman1966dynamic]. Please refer to the work in [zhu2025weakly] or our supplementary file for more details.
@@
VASNet [fajtl2018summarizing] 0.0712 0.0645 0.0676 100.62 0.43 0.0496 0.0422 0.0455 84.79 0.44
Video CLIP-It [narasimhan2021clip] 0.0711 0.0629 0.0667 101.12 0.45 0.0832 0.0710 0.0764 85.82 0.38
Summarization OTVS [wang2023self] 0.0688 0.0613 0.0648 101.25 0.46 0.0834 0.0711 0.0766 84.82 0.39
Muvee [ganhor2014muvee] 0.2400 0.0461 0.0714 103.50 0.36
M2T [rehusevych2019movie2trailer] 0.0611 0.0503 0.0515 95.67 0.42
V2T [irie2010automatic] 0.1121 0.0603 0.0945 103.75 0.52
Selection- PPBVAM [xu2015trailer] 0.0813 0.1244 0.0945 101.50 0.53
then-Ranking IPOT [wang2024inverse] 0.1218 0.1425 0.1311 101.50 0.42 0.1258 0.1483 0.1357 83.21 0.46
MMSC [zhu2025weakly] 0.1301 0.1449 0.1391 99.25 0.58 0.1852 0.2164 0.1991 82.48 0.50
Autoregression TGT [argaw2024towards] 0.1185 0.1142 0.1153 103.87 0.48 0.1243 0.1433 0.1326 88.29 0.43
@@ 4.1 Implementation Details
Datasets. We extend the CMTD dataset [wang2024inverse] to train our model and evaluate the model on the two testing sets of the dataset, called Test-8 and Test-74, respectively. This dataset contains 500 movies and 922 corresponding trailers. We enrich the dataset by 200 additional recent movies and their corresponding trailers. Among them, 30 movie-trailer pairs newly released in 2024 are selected to construct a new test set, called Test-2024, which is used to further evaluate the model’s generalization ability. The collection of all the three testing sets is called Test-ALL, which is applied in ablation studies. For each video (a movie or a trailer), we segment shots and extract their features following the pre-processing pipeline in [wang2024inverse, sun2025multi].
Baselines. We compare our SSMP method with state-of-the-art trailer generation methods, including selection-then-ranking methods (e.g., V2T [irie2010automatic], M2T [rehusevych2019movie2trailer], PPBVAM [xu2015trailer], IPOT [wang2024inverse], and MMSC [zhu2025weakly]) and the autoregressive method TGT [^2] Since the data and code are not publicly available, we reproduce TGT [argaw2024towards] by ourselves. When generating a trailer, we set the same length for TGT and our method for a fair comparison. [argaw2024towards, jin2025tvmtrailer]. Here, IPOT [wang2024inverse] and MMSC [zhu2025weakly] require background music as input, while the others only leverage visual information. Several video summarization methods are considered when evaluating the shot selection performance, including Muvee [ganhor2014muvee], VASNet [fajtl2018summarizing], CLIP-It [narasimhan2021clip], and OTVS [wang2023self]. For a fair comparison, all the methods are tested on the same dataset, and all the trailer generation methods apply the post-processing shown in Sec. 3.4.
Evaluation metrics. To evaluate the shot selection performance, we employ Precision, Recall, and F1-score (F1) by comparing each $\bm{z}$ with the ground-truth movie–trailer alignment. Following the work in [liu2023video, argaw2024towards, jin2025tvmtrailer], we further use Levenshtein Distance (LD) and Pairwise Agreement Accuracy (AA) to measure the shot ranking performance. For each generated trailer and the official one, LD quantifies their dissimilarity by computing the minimum number of edits required to convert one shot sequence into another, and AA measures the proportion of well-aligned shot pairs.
@@
VASNet [fajtl2018summarizing] 0.0990 0.0806 0.0886 93.84 0.44
CLIP-It [narasimhan2021clip] 0.0443 0.0370 0.0403 95.51 0.31
OTVS [wang2023self] 0.1021 0.0831 0.0914 93.56 0.44
IPOT [wang2024inverse] 0.1236 0.1460 0.1336 93.82 0.42
MMSC [zhu2025weakly] 0.1567 0.1485 0.1506 93.91 0.52
TGT [argaw2024towards] 0.1521 0.1601 0.1601 103.25 0.46